OpenStudy (anonymous):

Find y '' by implicit diferentiation sqrt of x + sqrt of y = 1

OpenStudy (anonymous):

\[\sqrt{x}+\sqrt{y} = 1\]

OpenStudy (anonymous):

\[\frac{1}{2} x^{-\frac{1}{2}} + \frac{1}{2} y^{- \frac{1}{2}} \frac{dy}{dx} = 0 \]

OpenStudy (anonymous):

\[\frac{1}{2 \sqrt{x}} + \frac{1}{2\sqrt{y}} \frac{dy}{dx} =0 \]

OpenStudy (anonymous):

\[\frac{dy}{dx} = - \sqrt{ \frac{y}{x}}\]

OpenStudy (anonymous):

y do u have to multiply that by dy/dx

OpenStudy (anonymous):

\[\frac{d}{dx} [ \frac{dy}{dx} ] = - \frac{d}{dx} [ (\frac{y}{x})^{\frac{1}{2}}] \]

OpenStudy (anonymous):

because it's a chain rule.

OpenStudy (anonymous):

were u isolating dy/dx to get -sqrt y/x?

OpenStudy (kira_yamato):

Yeah

OpenStudy (anonymous):

can u plz show the steps to how u got taht....

OpenStudy (anonymous):

i get confused if steps r skipped

OpenStudy (anonymous):

@kira u have dy/dx is positive whereas elec has it negaitve

OpenStudy (kira_yamato):

Oops... I left out -ve sign when I was typing for the screenshot... Gomen...

OpenStudy (anonymous):

im sorry but the math looks realllllly confusing...esp the last step where u were solving d^2y/dx^2

OpenStudy (anonymous):

do u mind explaing wat u were tryin to do

OpenStudy (kira_yamato):

Here's an errata, hope it doesn't contain any mistakes. I'm using quotient rule